Equilibrium was established for the reaction
2 HCl (g) H2 (g) + Cl 2(g)

At 2800 OC from an initial pressure of HCl of 0.35 atm. The equilibrium pressure of Cl2
was 0.022 atm.
a) What is the value of Kp for this equilibrium
1
Expert's answer
2020-12-06T13:56:28-0500

2HCl = H2 + Cl2

I 0.35 0 0

C -2x x x

E 0.35-2x x x


2x = pressure reduced from initial pressure from HCl.

So, now x = 0.022 ; pCl2 = 0.022

pH2 = 0.022 , pHCl = 0.35 - (2×0.022) =0.306 .

Kp = (pCl2×pH2) \ (pHCl)2

=(0.022×0.022) \ (0.306)2

Kp =5.17 × 10-3